OpenStudy (anonymous):

The circumference of C is 72 cm. what is the length of arc AB (the minor arc)? Step by step would be nice!

OpenStudy (anonymous):

Do you have a diagram at all?

OpenStudy (anonymous):

Here :)

OpenStudy (anonymous):

The formula for this is \[\frac{ n }{ 360 } \times 2\pi r\]

OpenStudy (anonymous):

But first you have to find the radius to use that.

OpenStudy (anonymous):

Oh okay I see, I was missing the n/360 part

OpenStudy (anonymous):

The formula for that is \[\frac{ c }{ 2\pi }\]

OpenStudy (anonymous):

Can you find it now?

OpenStudy (anonymous):

Yes. Thanks!

OpenStudy (anonymous):

You're welcome!